"What Is The X-Intercept For This Linear Equation? 6x-y=9"
Answer by stanbon(75887)

"What Is The X-Intercept For This Linear Equation? 6x-y=9"
---------------
Let y = 0,
then 6x-0 = 9
6x = 9
x = 9/6 = 3/2
===================
Cheers,
Stan H.
==========